OpenStudy (anonymous):

solve for the higher order derivative. find "y" if y = x^3 - 3x. please help me to solve this question.

OpenStudy (freckles):

"y" ?

OpenStudy (freckles):

does that mean second order derivative?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes sir

OpenStudy (freckles):

so you mean y''

OpenStudy (freckles):

first do you know how to find y'?

OpenStudy (freckles):

the power rule and constant multiple rule could come in handy here

OpenStudy (xapproachesinfinity):

i was wondering what "y" might mean lol

OpenStudy (anonymous):

i dont knw how to find y

OpenStudy (freckles):

y is given

OpenStudy (freckles):

we are trying to find y'

OpenStudy (xapproachesinfinity):

you are doing derivative of x^3 and -3x

OpenStudy (freckles):

\[(x^n)'=nx^{n-1} \text{ is power rule }\]

OpenStudy (anonymous):

i think combine terms

OpenStudy (freckles):

\[(cx)' =c(x)' \text{ by constant multiple rule } \\ =c(1) \text{ by the rule } \frac{dx}{dx}=1 \\ =c \text{ since 1 is multiplicative identity } \]

OpenStudy (freckles):

there are no like terms to combine

OpenStudy (anonymous):

then?

OpenStudy (freckles):

have you ever seen the rules I have mentioned?

OpenStudy (anonymous):

how can we solve that problem?

OpenStudy (xapproachesinfinity):

you need to know those rules to do this!!!

OpenStudy (freckles):

If you don't know these rules we can use the definition of derivative

OpenStudy (freckles):

but i need you to tell me if you do or do not know them

OpenStudy (anonymous):

i do not know them

OpenStudy (anonymous):

sorry

OpenStudy (freckles):

\[\frac{df(x)}{dx}=\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h}\] have you seen this definition before?

OpenStudy (anonymous):

no

OpenStudy (freckles):

hmm... we are out out of luck then

OpenStudy (freckles):

because without knowing those rules and/or that definition there is no way I see how we can find the derivative

OpenStudy (anonymous):

but i have a formula that ive know

OpenStudy (freckles):

oh what is it

OpenStudy (anonymous):



OpenStudy (anonymous):

y^1= x^3 - 3x

OpenStudy (anonymous):

it will gonna be like this right?

OpenStudy (freckles):

that is the quotient rule

OpenStudy (freckles):

we don't have a quotient

OpenStudy (freckles):

how do you know about the quotient rule but not anything about the constant multiple rule, constant rule, or the power rule or even the definition of derivative?

OpenStudy (freckles):